OpenStudy (anonymous):

add and simplify 4/9 + 1/2

Parth (parthkohli):

= 8/18 + 9/18 = 17/18

OpenStudy (anonymous):

\[= 8/18 + 9/18\] \[\rightarrow 17/18\]]